# GraspXL: Generating Grasping Motions for Diverse Objects at Scale

### v1

@@ 4.2 Method Comparison
PartNet Evaluation: We use the PartNet test set to evaluate all methods with the given objectives. For each object, we calculate the average performance among 25 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$. Table 2 shows that our method significantly outperforms the baselines across all metrics. In particular, our method more closely follows the objectives in terms of hand grasping position (Midpoint Error), heading direction (Heading Error), wrist rotation (Wrist Rotation Error), and contact points with the graspable/non-graspable parts (Contact Ratio). At the same time, we achieve the most stable grasps (Success Rate).
ShapeNet Evaluation: While PartNet specializes in part-based objects at a smaller scale, we extend our evaluation to ShapeNet to test on more diverse and unseen objects. We measure average performance across 5 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$ per object. As shown in Table 2, our method outperforms all baselines, demonstrating less decrease in performance when comparing PartNet and ShapeNet, thereby proving superior generalization. The Grasp Success Rate has a drop for all the methods (even for the non-learning-based SynH2R-PD baseline), which we claim is caused by some objects that are too large or too heavy to be grasped. It is worth noting that generating grasp pose references for the ShapeNet test set using the baselines requires approximately a week, whereas our method can directly perform real-time inference in simulation, showcasing its efficiency. Please refer to SupMat for further qualitative comparisons.
@@ 4.3 Generalization
Generalization to Large-scale Object Dataset: Our method’s scalability was tested using the large Objaverse [12] dataset, resulting in a test set of over half a million objects of three different sizes: small, medium and large (refer to SupMat for preprocessing details). Performance on this set, as detailed in Table 3, is comparable to our ShapeNet results, underscoring our method’s ability to scale. Specifically, medium-sized objects offer optimal grasping success. Smaller objects enable higher success rates and more accurate grasp positioning but suffer from larger heading direction and wrist rotation errors, as smaller objects are easier to grasp but also easier to rotate in hand. Overall, our method consistently shows excellent performance across different object sizes, confirming its effectiveness for varying scales.
@@
Generalization to Reconstructed and Generated Objects: Our method not only synthesizes grasping motions for standard 3D assets [12, 29, 4] but also effectively handles reconstructed and generated objects, expanding its usability. We evaluated its performance on objects reconstructed via HOLD [17] and those generated by DreamFusion [30]. We analyzed the average performance for each object across 25 random sets of motion objectives, with findings detailed in Tab. 4. The results are presented in Tab. 4. Despite the objects being novel and containing severe artifacts (see the Rubik’s Cube in Fig. 2), our method maintains similar performance across all metrics compared to other experiments. Notably, the performance of reconstructed HO3D objects is similar with the performance on ground-truth objects. This shows our method’s robustness towards reconstruction noise in object meshes. Please refer to the SupMat for a more detailed explanation of the setting.
@@
Figure 5: Generated Motions of Different Hands with the Same Objectives. We require the hands to approach from right and grasp the upper part of the glass.
Generalization on Different Robotic Hands: We assess our framework’s generalization capabilities on various dexterous robotic hands, including Shadow Hand [36], Allegro Hand [44], and Faive Hand [41], each differing in size and joint structure (Faive Hand has 30, Shadow Hand 22, Allegro Hand 16, and MANO 45 finger joints). Our method adapts seamlessly to different models by adjusting hand state and action space dimensions. According to Tab. 5, all tested hands showed similar success rates. However, the Allegro Hand has greater Mid. Errors and lower Contact Ratios due to its larger size, which makes it difficult to achieve precise position control. The Shadow Hand shows marginally higher objective errors due to its broad, flat palm which limits its dexterity. Despite these structural differences affecting performance metrics, all hands achieved commendable results, highlighting our method’s ability to generalize. This suggests our framework’s capacity to deal with hand models with different morphologies. Figure 5 illustrates the variety of grasping motions generated for different hands under the same motion objectives.
@@ 4.4 Ablation and Analysis
Ablation: We conduct an ablation study to evaluate the impact of different components in our method. Specifically, we assessed variations without the hand guidance technique (w/o Guidance), the joint distance features (w/o Distance), and the learning curriculum (w/o Curriculum). Results are detailed in Table 6. The version without the curriculum slightly outperforms in contact ratio on the PartNet set and maintains similar success rates across tests but suffers significantly in objective-related metrics. We also notice the absence of the curriculum makes the results more sensitive to the exact reward coefficients, leading to more grasp failure with slightly higher objective rewards or poor objective precision with higher grasping rewards. These underscore the curriculum’s role in decoupling the learning of stable grasping and objective fulfillment, which can help the policy avoid getting stuck in local optima caused by their influence on each other during training. The model without the joint distance features gets a slightly larger midpoint error and significantly underperforms in all other aspects, highlighting the distance features’ value in understanding object shapes and adjusting to contact-induced movements. Lastly, excluding hand guidance worsens all performance metrics, proving its efficacy in directing the hand to achieve the desired grasp.
@@
Evaluation on Different Objective Combinations: Our method’s adaptability was tested across various objective combinations to gauge performance impacts. Evaluations were conducted for different scenarios: solely controlling heading direction ($\bar{\mathbf{m}}$), adding wrist rotation to heading direction ($\bar{\mathbf{v}}+\bar{\omega}$), combining heading direction with midpoint position ($\bar{\mathbf{v}}+\bar{\mathbf{m}}$), and integrating all three objectives ($\bar{\mathbf{v}}+\bar{\omega}+\bar{\mathbf{m}}$). This variety helps to understand our approach’s efficacy in handling multiple, concurrent objectives and understanding the complexity levels our method can effectively manage. For each objective set, we sample 25 sets of motion objective for PartNet objects and 5 for ShapeNet objects. Results in Tab. 7 show slightly better performance with fewer motion objectives, highlighting the increased control challenge with multiple objectives. This demonstrates our method’s capability to handle varying difficulty levels and its robustness, as indicated by the minimal performance decline.
@@ 5 Conclusion
The supplementary material contains this document and a video. In Section 6, we provide qualitative results compared against our baseline. We then provide the implementation details about hyperparameters and motion objectives in Section 7. In Section 8, we show the experiment details about data preprocessing and the evaluation with generated and reconstructed objects. Finally, we provide additional experiments in Section 9. We will release code and models.
@@ 6 Qualitative Results
We provide qualitative comparisons of our method with SynH2R [7] for objective-driven grasping synthesis in Figure 6. From the figures, we can see that SynH2R either failed to establish a stable grasp due to noisy generated references, or cannot precisely follow the objectives. However, our method does not require a pre-generated reference, and can generate motions with stable grasping while satisfying motion objectives.
@@ 7.1 Training Hyperparameters
We use PPO [35] to train our policy and follow the implementation provided in [9]. We present an overview of the important parameters and weight values of the reward function in Table 8 and Table 9.
@@ 9.1 Training Set Size Effect Evaluation
To show the data efficiency of our method, we train another policy with an enlarged training set composed of 100 PartNet objects (with the graspable area partitions), and 400 ShapeNet objects. We then perform the same evaluation with the PartNet and ShapeNet test sets (See Section 4.1 in the main manuscript). The results are shown in Table 10. Compared with the results with a smaller training set (See Section 4.2 in the main manuscript), there is no significant improvement, which means that a training set composed of 58 objects with diverse shape is sufficient for our framework. This shows the data efficiency of our method. We hypothesize that this is because diverse shapes together with random objectives and initialization during training provide a diverse distribution of grasps and configurations.
@@
9.2 SOTA Comparison for Robotic Hand Grasping
To show the effectiveness of our method on robotic applications, we compare our method with two state-of-the-art dexterous robot hand manipulation methods, UniDexGrasp [45] and UniDexGrasp++ [43]. UniDexGrasp samples and optimizes diverse grasp pose references, filters the successful ones with a physics simulator, and then generates diverse grasp motions accordingly. UniDexGrasp++ is specifically trained for higher grasp success rates and generates a single pose for each object. We evaluate our method using their hand model (the Shadow Hand), test sets (one for seen categories and one for unseen categories), and metric (success rate), and compare with their state-based policies. The results are shown in Table 11. Note that both their test sets are unseen to our policy. Our policy is trained with only our 58 training objects instead of their 3000+ training objects. For each object, we randomly sample 25 target approach directions to synthesize diverse grasp motions. Our method outperforms both methods on both test sets. Specifically, although UniDexGrasp only uses filtered references for motion synthesis while ours can directly inference in simulation, our method significantly outperforms it and retains the ability to generate diverse motions. Notably, our method achieves a higher success rate than UniDexGrasp++ on its seen category test set.
Table 11: Comparison with Dexterous Robot Hand Manipulation Methods. We evaluate the performance of our policy using the task setting and metric of UniDexGrasp and UniDexGrasp++ on their test sets. Both ours and UniDexGrasp generate diverse grasps, while UniDexGrasp++ only considers a single grasp for each object.
Method Success Rate [%] $\uparrow$ Success Rate [%] $\uparrow$
(Seen Category) (Unseen Category)
UniDexGrasp (Diverse poses) 71 67
UniDexGrasp++ (Single pose) 84 83
Ours (Diverse poses) 85 89

### v2

@@ 4.2 Method Comparison
PartNet Evaluation: We use the PartNet test set to evaluate all methods with the given objectives. For each object, we calculate the average performance among 25 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$. Tab. 2 shows that our method significantly outperforms the baselines across all metrics. In particular, our method more closely follows the objectives in terms of hand grasping position (Midpoint Error), heading direction (Heading Error), wrist rotation (Wrist Rotation Error), and contact points with the graspable/non-graspable parts (Contact Ratio). At the same time, we achieve the most stable grasps (Grasping Success Rate).
ShapeNet Evaluation: While PartNet specializes in part-based objects at a smaller scale, we extend our evaluation to ShapeNet to test on more diverse and unseen objects. We measure average performance across 5 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$ per object. As shown in Tab. 2, our method outperforms all baselines, demonstrating less decrease in performance when comparing PartNet and ShapeNet, thereby proving superior generalization. The Grasping Success Rate has a drop for all the methods (even for the non-learning-based SynH2R-PD baseline), which we claim is caused by some objects that are too large or too heavy to be grasped. It is worth noting that generating grasp pose references for the ShapeNet test set using the baselines requires approximately a week, whereas our method can directly perform real-time inference in simulation, showcasing its efficiency. Please refer to SupMat for further qualitative comparisons.
@@ 4.3 Generalization
Generalization to Large-scale Object Dataset: Our method’s scalability was tested using the large Objaverse [12] dataset, resulting in a test set of over half a million objects of three different sizes: small, medium and large (refer to SupMat for preprocessing details). Performance on this set, as detailed in Tab. 3, is comparable to our ShapeNet results, underscoring our method’s ability to scale. Specifically, medium-sized objects offer optimal grasping success. Smaller objects enable higher success rates and more accurate grasp positioning but suffer from larger heading direction and wrist rotation errors, as smaller objects are easier to grasp but also easier to rotate in hand. Overall, our method consistently shows excellent performance across different object sizes, confirming its effectiveness for varying scales. The generated grasping motions of different hands are released for further research.
@@
Generalization to Reconstructed and Generated Objects: Our method not only synthesizes grasping motions for standard 3D assets [12, 30, 4] but also effectively handles reconstructed and generated objects, expanding its usability. We evaluated its performance on objects reconstructed via HOLD [17] and objects generated by DreamFusion [31]. We analyzed the average performance for each object across 25 random sets of motion objectives, with findings detailed in Tab. 4. Despite the objects being novel and containing severe artifacts (see the Rubik’s Cube in Fig. 2), our method maintains similar performance across all metrics compared to other experiments. Notably, the performance of objects reconstructed from HO3D [21] is similar with the performance of ground-truth objects. This shows our method’s robustness towards reconstruction noise in object meshes. Please refer to the SupMat for a more detailed setting explanation.
@@
Figure 5: Generated Motions of Different Hands with the Same Objectives. We require the hands to approach from the right and grasp the upper part of the glass.
Generalization on Different Robotic Hands: We assess our framework’s generalization capabilities on various dexterous robotic hands, including Shadow Hand [37], Allegro Hand [44], and Faive Hand [41], each differing in size and joint structure (Faive Hand has 30, Shadow Hand 22, Allegro Hand 16, and MANO 45 finger joints). Our method adapts seamlessly to different models by adjusting hand state and action space dimensions. According to Tab. 5, all tested hands showed similar success rates. However, the Allegro Hand has greater Mid. Errors and lower Contact Ratios due to its larger size, which makes it difficult to achieve precise position control. The Shadow Hand shows marginally higher objective errors due to its broad, flat palm which limits its dexterity. Despite these structural differences affecting performance metrics, all hands achieved commendable results, highlighting our method’s ability to generalize. This suggests our framework’s capacity to deal with hand models with different morphologies. Fig. 5 illustrates the variety of grasping motions generated for different hands under the same motion objectives.
@@ 4.4 Ablation and Analysis
Ablation: We conduct an ablation study to evaluate the impact of different components in our method. Specifically, we assessed variations without the hand guidance technique (w/o Guidance), the joint distance features (w/o Distance), and the learning curriculum (w/o Curriculum). Results are detailed in Tab. 6. The version without the curriculum slightly outperforms in contact ratio on the PartNet set and maintains similar success rates across tests but suffers significantly in objective-related metrics. We also notice the absence of the curriculum makes the results more sensitive to the exact reward coefficients, leading to more grasp failure with slightly higher objective rewards or poor objective precision with higher grasping rewards. These underscore the curriculum’s role in decoupling the learning of stable grasping and objective fulfillment, which can help the policy avoid getting stuck in local optima caused by their influence on each other during training. The model without the joint distance features gets a slightly larger midpoint error and significantly underperforms in all other aspects, highlighting the distance features’ value in understanding object shapes and adjusting to contact-induced movements. Lastly, excluding hand guidance worsens all performance metrics, proving its efficacy in directing the hand to achieve the desired grasp.
@@
Evaluation on Different Objective Combinations: Our method’s adaptability was tested across various objective combinations to gauge performance impacts. Evaluations were conducted for different scenarios: solely controlling heading direction ($\bar{\mathbf{m}}$), adding wrist rotation to heading direction ($\bar{\mathbf{v}}+\bar{\omega}$), combining heading direction with midpoint position ($\bar{\mathbf{v}}+\bar{\mathbf{m}}$), and integrating all three objectives ($\bar{\mathbf{v}}+\bar{\omega}+\bar{\mathbf{m}}$). This variety helps to understand our approach’s efficacy in handling multiple, concurrent objectives and understanding the complexity levels our method can effectively manage. For each objective set, we sample 25 sets of motion objectives for PartNet objects and 5 for ShapeNet objects. Results in Tab. 7 show slightly better performance with fewer motion objectives, highlighting the increased control challenge with multiple objectives. This demonstrates our method’s capability to handle varying difficulty levels and its robustness, as indicated by the minimal performance decline.
@@ 5 Conclusion
In Sec. 6, we provide qualitative results compared against our baseline. We then provide the implementation details about hyperparameters and motion objectives in Sec. 7. In Sec. 8, we show the experiment details about data preprocessing and the evaluation with generated and reconstructed objects. Finally, we provide additional experiments in Sec. 9. Our model, code, and the large-scale generated motions are released for future research. Check our project page: https://eth-ait.github.io/graspxl/ for more details and visualization.
@@ 6 Qualitative Results
We provide qualitative comparisons of our method with SynH2R [8] for objective-driven grasping synthesis in Fig. 6. From the figures, we can see that SynH2R either failed to establish a stable grasp due to noisy generated references, or cannot precisely follow the objectives. However, our method does not require a pre-generated reference, and can generate motions with stable grasping while satisfying motion objectives.
@@ 7.1 Training Hyperparameters
We use PPO [36] to train our policy and follow the implementation provided in [9]. We present an overview of the important parameters and weight values of the reward function in Tab. 8 and Tab. 9.
@@ 8.1 Dataset Preprocessing
As the object meshes from the datasets contain no material information for density and friction, we calculate object masses based on the mesh volume for a given fixed density, leading to diverse masses. We use the same friction coefficient in simulation for all objects.
@@ 9.1 Training Set Size Effect Evaluation
To show the data efficiency of our method, we train another policy with an enlarged training set composed of 100 PartNet objects (with the graspable area partitions), and 400 ShapeNet objects. We then perform the same evaluation with the PartNet and ShapeNet test sets (See Section 4.1 in the main manuscript). The results are shown in Tab. 10. Compared with the results with a smaller training set (See Section 4.2 in the main manuscript), there is no significant improvement, which means that a training set composed of 58 objects with diverse shape is sufficient for our framework. This shows the data efficiency of our method. We hypothesize that this is because diverse shapes together with random objectives and initialization during training provide a diverse distribution of grasps and configurations.
@@
9.2 Friction Effect Evaluation
Although we use the same friction coefficient in simulation for all objects, our method can also deal with randomized frictions (±0.3 around the default friction coefficient) as shown in Tab. 11.
Table 11: Evaluation with random friction coefficients.
Settings Suc. Rate [%] $\uparrow$ Mid. Error $\downarrow$
Original experiment 81.6 0.283
Random friction coefficient 80.5 0.285
9.3 Realism Evaluation
To evaluate the realism of our method, we invite 35 participants to score the realism in terms of human likeness, naturalness, smoothness, and hand-object interpenetration. They score 10 sets of rendered grasping motions from 1 (worst) to 3 (best), each containing three motions of the same object randomly chosen from ours, HO3D [21], and DexYCB [5]. The average scores are 2.12, 2.05, and 2.45, respectively. Our method has a slightly higher realism score than HO3D as HO3D has some interpenetrations and jitters caused by labeling noise, and DexYCB has the highest score due to accurate annotations.
9.4 Inference Speed Evaluation
With a 24-core Intel i9-14900 CPU and an Nvidia RTX 4090 GPU, we generate 100 sequences and calculate the average time consumption per sequence with our method, the baselines SynH2R-PD, and SynH2R [8], leading to 37.15, 37.17, and 0.14 seconds, respectively. The most time consumption for SynH2R-PD and SynH2R are caused by the optimization-based static grasping reference pose generation procedure.